If f(x)=x^2-5x+8 & g(x)=x^2-4, which is an equivalent form of g(x)+f(x)?
A.x^2-5x+4
B.2x^2-5x-4
C.2x^2-5x+4
D.2x^2+5x+4
Your Help We Be Appreciated.

Answer by jim_thompson5910(35256)

g(x)+f(x) = x^2-4+x^2-5x+8 = 2x^2-5x+4